Marivic is 20 years older than Fe. Seven years from now , Marivic will be thrice as old as Fe. Find Their present ages
Answer by ankor@dixie-net.com(15645)

Write an equation for each statmente
:
Marivic is 20 years older than Fe.
m = f + 20
:
Seven years from now , Marivic will be thrice as old as Fe.
m + 7 = 3(f+7)
m + 7 = 3f + 21
m = 3f + 21 - 7
m = 3f + 14
Substitute (f+20) from the 1st statement:
f + 20 = 3f + 14
20 - 14 = 3f - f
6 = 2f
f = 3 yrs
and
m = 23 yrs
:
:
Check solution in the statement:
"Seven years from now , Marivic will be thrice as old as Fe. "
23 + 7 = 3(3 + 7)